# Activation of entanglement in generalized entanglement swapping

### v2

@@ Abstract
We study entanglement activation in a generalized entanglement swapping process involving two Bell pairs and generalized measurements. The conventional understanding posits entangled measurements as both necessary and sufficient for establishing entanglement between distant parties. In this study, we reassess the role of measurement operators in entanglement generation within a generalized entanglement swapping process. We focus on maximally entangled two-qubit initial states and generalized measurements, investigating the necessity and sufficiency conditions for entangled measurement operators. By utilizing two Bell pairs, (1, 2) shared between Alice and Bob, and (3, 4) shared between Bob and Charlie, we demonstrate that while entangled measurements are sufficient, they are not indispensable for establishing entanglement between spatially separated observers. Through a sequential approach, if Bob performs an initial measurement which is not able to establish entanglement then followed by another measurement after post-processing the first measurement it is possible to establish entanglement. We identify specific criteria for different measurement operators that enable the potential for performing a second measurement to establish entanglement. Our findings highlight the feasibility of generating entanglement between distant parties through a combination of measurements, shedding new light on entanglement distribution in quantum networks. Additionally, we showcase through illustrative examples how successive measurements enhance entanglement compared to single measurements, underscoring the practical benefits of our approach in enhancing entanglement. Moreover, our protocol extends beyond bipartite qubit states to higher-dimensional maximally entangled states, emphasizing its versatility and applicability.
@@ I INTRODUCTION
In operational terms, an operator is referred to as an inseparable operator if its Choi state is entangled while separable operators are operators which take separable states to separable states and its Choi state is a separable density matrix rain1 ; rain2 ; rain3 . Quantum measurements are a special class of operators. The most general form of measurement is called a Positive Operator-Valued Measure (POVM) um1 ; unsharp_measurement ; povm . A POVM, $\mathcal{M}$ is defined as a set of positive operators $\Pi_{n}$ s.t $\sum_{n}{\Pi_{n}}=\mathbb{I}$ where $\mathbb{I}$ is the identity operator on the corresponding Hilbert space. Measurements are characterized by various definitions. In papers yokoyama ; hamamura ; sperling ; duan ; vertesi ; rabelo ; adam ; wu ; virmani , it has been defined that a POVM element $\Pi_{n}$ is considered entangled if the partial transpose of the corresponding operator, defined as $\frac{\Pi_{n}}{\mbox{Tr}(\Pi_{n})}$, fails to be positive. The POVM element is said to be unentangled if the normalized form $\frac{\Pi_{n}}{\mbox{Tr}(\Pi_{n})}$ is a separable quantum mixed state. A POVM measurement is called an entangled measurement if at least one of the POVM elements is entangled. Similarly, A POVM measurement is called unentangled measurement if each of the POVM elements is unentangled. Here, the entanglement of the measurement is characterized as an entanglement of the detection devices or POVM entanglement. The entire measurement set is classified into different categories based on this definition vertesi ; adam . In vertesi ; adam , the authors have classified the set of measurements into different categories such as entangled measurement, unentangled measurement, LOCC measurement, and classical measurement where classical and LOCC measurements do not have any capacity to produce the entangled state from a separable state. Note that these measurements have the inclusion relation: General measurement $\supset$ unentangled measurement $\supset$ LOCC measurement $\supset$ classical measurement, where general measurement corresponds to the set of all quantum measurements, including the entangled measurements vertesi ; adam . LOCC and classical measurements belong to the class of separable operations in an operational sense. It is important to note that, according to this definition, not every unentangled measurement is considered a separable operation, but the converse is true.
@@ II GENERALIZED PROTOCOL
$|\Phi\rangle=|\phi^{+}\rangle_{12}\otimes|\phi^{+}\rangle_{34}$ (1)
@@
where $|\phi_{n\alpha}\rangle$, $\alpha=1,2,3,4$ form a complete orthonormal basis on $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, $\pi_{n\alpha}\geq 0$ for $\alpha=1,2,3,4$. Here, $K$ represents the number of POVM elements in a POVM measurement. If Bob performs a joint two-qubit generalized quantum measurement $\mathcal{M}_{1}$ on ($2,3$) of the $|\Phi\rangle$ state then the post-measurement state of the $n$-th outcome is
@@
with probability $p_{n}$ where it is understood that $\Pi_{n}$ is acting on (2,3) pair and identity operator ($\mathbb{I}_{4}$) acting on (1,4) pair. The probability of obtaining $n$-th outcome is
@@
$=$ $\frac{1}{4}\mbox{Tr}(\Pi_{n})$
Using Eq.1 and Eq.2 in Eq.3 pbej1 , one can get
$|\Phi_{n}\rangle=\frac{1}{2\sqrt{p_{n}}}\sum_{\alpha=1}^{4}\sqrt{\pi_{n\alpha}}|\phi_{n\alpha}\rangle_{23}|\phi_{n\alpha}^{*}\rangle_{14}$ (5)
where $|\phi^{*}\rangle$ denotes the complex conjugate of $|\phi\rangle$ in the computational basis. Note that after the measurement, the total system state Eq.5 remains pure.
Now after tracing out (2,3) qubits the post-measurement state between Alice and Charlie for $n$-th outcome becomes
@@
$=$ $\frac{\Pi_{n}^{*}}{\mbox{Tr}(\Pi_{n})}$
where $\Pi_{n}^{*}=\sum_{\alpha=1}^{4}\pi_{n\alpha}|\phi_{n\alpha}^{*}\rangle\langle\phi_{n\alpha}^{*}|$. We also determine the post-measurement state of (1,2) and (3,4) pairs; for details, see Appendix A. Equation 6 illustrates how Bob’s measurement fully determines the state of the pair (1, 4). This equation also highlights the direct relationship between the post-measurement state of the (1,4) pair and the POVM measurement element. Conventional thought suggests that entanglement at the (1,4) pair requires an entangled measurement. However, we demonstrate that although an entangled POVM is sufficient to obtain an entangled state at (1,4), it is not a necessary condition
@@
Now, we consider that Bob post-processes each outcome of the first measurement $\mathcal{M}_{1}$ and then performs a second measurement $\mathcal{M}_{2}$ on the (2,3) qubits of the total system state Eq.5. Additionally, Bob classically communicates the outcomes to Alice and Charlie during the measurement process.
@@
$E_{m}=\sum_{\delta=1}^{4}\mu_{m\delta}|\psi_{m\delta}\rangle\langle\psi_{m\delta}|$ (7)
where $|\psi_{m\delta}\rangle$, $\delta=1,2,3,4$ form a complete orthonormal basis on $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, $\mu_{m\delta}\geq 0$ for $\delta=1,2,3,4$.
The post-measurement state of the four-qubit system, when Bob’s outcomes for the successive measurements are the $n$-th element of the measurement $\mathcal{M}_{1}$ and the $m$-th element of $\mathcal{M}_{2}$ is given by:
@@
where it is understood that $E_{m}$ is acting on (2,3) pair and $\mathbb{I}_{4}$ is acting on (1,4) pair of $|\Phi_{n}\rangle$ which is given by Eq.5. The probability of obtaining the $|\Phi_{nm}\rangle$ as the outcome state is $p_{n}s_{nm}$.
Here, we denote by $s_{nm}$ the probability of obtaining the $m$-th outcome of the second measurement, $\mathcal{M}_{2}$ when it acts on the $n$-th outcome of the first measurement $\mathcal{M}_{1}$. Using Eq.4, Eq.5, and Eq.7, we get
@@
Now using Eq.5, and Eq.7, the above Eq.8 can be written as
$|\Phi_{nm}\rangle$ $=$ $\frac{1}{2\sqrt{p_{n}s_{nm}}}\sum_{\alpha,\delta=1}^{4}\sqrt{\pi_{n\alpha}}\sqrt{\mu_{m\delta}}\langle\psi_{m\delta}|\phi_{n\alpha}\rangle|\psi_{m\delta}\rangle_{23}\otimes|\phi_{n\alpha}\rangle_{14}$
where $p_{n}$ and $s_{nm}$ are given by Eq.4 and Eq.9 respectively.
After tracing out the (2,3) pair from the above Eq.II, we obtain the post-measurement state between A and C for each outcome, given by
$\rho_{14|nm}=\frac{1}{4p_{n}s_{nm}}\sum_{\alpha,\beta=1}^{4}\sum_{\delta=1}^{4}\mu_{m\delta}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\langle\phi_{n\alpha}|\psi_{m\delta}\rangle\langle\psi_{m\delta}|\phi_{n\beta}\rangle|\phi_{n\alpha}\rangle\langle\phi_{n\beta}|$ (11)
where $|\phi_{n\alpha(\beta)}\rangle$, $\alpha(\beta)=1,2,3,4$ are orthonormal basis of the $n$-th element $\Pi_{n}$ of the first measurement $\mathcal{M}_{1}$.
@@
Lemma-1: If the first measurement $\mathcal{M}_{1}$ is a projective or rank-1 POVM (pure POVM) measurement, meaning each $\Pi_{n}$ element is rank-1 or projective, then any second measurement $\mathcal{M}_{2}$ cannot disturb the post-measurement state of $\mathcal{M}_{1}$ between Alice and Charlie, whether by entangling, disentangling, or increasing the entanglement.
Proof: A POVM, $\mathcal{M}$ is called as rank-1 when all its elements are rank-1 operators. Suppose $\mathcal{M}_{1}$ is a projective or rank-1 POVM, implying each element $\Pi_{n}$ of $\mathcal{M}_{1}$ measurement is rank-1, such that for $\alpha=1$, $\Pi_{n}=\pi_{n1}|\phi_{n1}\rangle\langle\phi_{n1}|$. Then, according to Eq.5, $|\Phi_{n}\rangle$ can be expressed as:
@@
Thus far, we establish that for the second measurement $\mathcal{M}_{2}$ to disturb the (1,4) pair in the post-measurement state $|\Phi_{n}\rangle$, two conditions must be met: Firstly, the rank of the first measurement $\mathcal{M}_{1}$ should exceed one, and secondly, $|\Phi_{n}\rangle$ should exhibit non-zero entanglement in the $14|23$ bipartition (Eq.14). This prompts the question: What are the properties of measurements that can lead to entangled states in (1,4) through successive measurements if the first attempt fails? To address this inquiry, we present protocols for different categories of measurements in the entanglement swapping scenario
@@ II.2 Separable operations
where we assume $|\Psi\rangle_{12}=|\phi^{+}\rangle$, a Bell state. Eq.30 reveals that any rank-2 measurement element retains the entanglement in the (1,2) and (3,4) pairs. Therefore, in the second round of measurement, if Bob conducts an inseparable operation on the (2,3) pair after post-processing the first measurement $\mathcal{M}_{1}$, there’s a possibility of obtaining an entangled state in the (1,4) pair. However, it’s crucial to note that the second measurement should be an inseparable operation, ensuring that the condition $\mathcal{C}_{n|12vs34}>0$ for an outcome (as stated in Eq. 17) is met.
It is noteworthy that our protocols are not only applicable for two rounds of successive measurements, but one can proceed further for n-rounds by following the same procedure until the (1,4) pair becomes entangled. After each round of measurement, the entanglement between the $14|23$ bipartition of the four qubits post-measurement state must be non-zero, for an outcome, in order for the next round of measurement to disturb the (1,4) pair, after post-processing the previous round measurements. We already know that a measurement will be successful in establishing entanglement in the (1,4) pair if it can create entanglement between the $12|34$ bipartition of the total system state.
@@ III CONCLUSIONS
Our findings demonstrate that while an entangled measurement is sufficient, it is not mandatory for the measurement operators to be entangled to establish an entangled state between A and C. We introduce a sequential approach where Bob conducts an initial measurement, followed by another measurement if the initial one fails to establish the entanglement. This sequential strategy, in conjunction with appropriate post-processing after the initial measurement, facilitates the establishment of entanglement between A and C. Our investigation offers new insights into the role of measurement operators in entanglement generation.
@@
Future research efforts should prioritize refining entanglement activation protocols, exploring innovative measurement strategies, and broadening the protocol’s applicability to diverse scenarios, including nonlocality activation klobus and network nonlocality bran1 ; bran2 . By addressing these research directions, we can deepen our understanding of entanglement generation and distribution, leading to advancements in quantum communication and computation.
@@ Acknowledgements
The authors express gratitude to Prof. Somshubhro Bandyopadhyay for fruitful discussions. PB acknowledges Saronath Halder for helpful discussions.

### v3

@@ Abstract
We study entanglement activation in a generalized entanglement swapping process involving two Bell pairs and generalized measurements. The conventional understanding posits entangled measurements as both necessary and sufficient for establishing entanglement between distant parties. In this study, we reassess the role of measurement operators in entanglement generation within a generalized entanglement swapping process. We focus on maximally entangled two-qubit initial states and generalized measurements, investigating the necessity and sufficiency conditions for entangled measurement operators. By utilizing two Bell pairs, (1, 2) shared between Alice and Bob, and (3, 4) shared between Bob and Charlie, we demonstrate that while entangled measurements are sufficient, they are not indispensable for establishing entanglement between spatially separated observers. Through a sequential approach, if Bob performs an initial measurement which is not able to establish entanglement then followed by another measurement after post-processing the first measurement it is possible to establish entanglement. We identify specific criteria for different measurement operators that enable the potential for performing a second measurement to establish entanglement. Our findings highlight the feasibility of generating entanglement between distant parties through a combination of measurements, shedding light on entanglement distribution in quantum networks. Additionally, we showcase through illustrative examples how successive measurements enhance entanglement compared to single measurements, underscoring the practical benefits of our approach in enhancing entanglement. Moreover, our protocol extends beyond bipartite qubit states to higher-dimensional maximally entangled states, emphasizing its versatility and applicability.
@@ I INTRODUCTION
In operational terms, an operator is referred to as an inseparable operator if its Choi state is entangled while separable operators are operators which take separable states to separable states and its Choi state is a separable density matrix rain1 ; rain2 ; rain3 . Quantum measurements are a special class of operators. The most general form of measurement is called a positive operator-valued measure (POVM) um1 ; unsharp_measurement ; povm . A POVM, $\mathcal{M}$ is defined as a set of positive operators $\Pi_{n}$ so that $\sum_{n}{\Pi_{n}}=\mathbb{I}$ where $\mathbb{I}$ is the identity operator on the corresponding Hilbert space. Measurements are characterized by various definitions. In papers yokoyama ; hamamura ; sperling ; duan ; vertesi ; rabelo ; adam ; wu ; virmani , it has been defined that a POVM element $\Pi_{n}$ is considered entangled if the partial transpose of the corresponding operator, defined as $\frac{\Pi_{n}}{\mbox{Tr}(\Pi_{n})}$, fails to be positive. The POVM element is said to be unentangled if the normalized form $\frac{\Pi_{n}}{\mbox{Tr}(\Pi_{n})}$ is a separable quantum mixed state. A POVM measurement is called an entangled measurement if at least one of the POVM elements is entangled. Similarly, A POVM measurement is called unentangled measurement if each of the POVM elements is unentangled. Here, the entanglement of the measurement is characterized as an entanglement of the detection devices or POVM entanglement. The entire measurement set is classified into different categories based on this definition vertesi ; adam . In vertesi ; adam , the authors have classified the set of measurements into different categories such as entangled measurement, unentangled measurement, local operations and classical communication (LOCC) measurement, and classical measurement where classical and LOCC measurements do not have any capacity to produce the entangled state from a separable state. Note that these measurements have the inclusion relation: General measurement $\supset$ unentangled measurement $\supset$ LOCC measurement $\supset$ classical measurement, where general measurement corresponds to the set of all quantum measurements, including the entangled measurements vertesi ; adam . LOCC and classical measurements belong to the class of separable operations in an operational sense. It is important to note that, according to this definition, not every unentangled measurement is considered a separable operation, but the converse is true.
@@ II GENERALIZED PROTOCOL
$|\Phi\rangle=|\phi^{+}\rangle_{12}\otimes|\phi^{+}\rangle_{34},$ (1)
@@
where $|\phi_{n\alpha}\rangle$, $\alpha=1,2,3,4,$ form a complete orthonormal basis on $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, $\pi_{n\alpha}\geq 0$ for $\alpha=1,2,3,4$. Here, $K$ represents the number of POVM elements in a POVM measurement. If Bob performs a joint two-qubit generalized quantum measurement $\mathcal{M}_{1}$ on ($2,3$) of the $|\Phi\rangle$ state, then the post-measurement state of the $n$th outcome is
@@
with probability $p_{n}$ where it is understood that $\Pi_{n}$ is acting on (2,3) pair and identity operator ($\mathbb{I}_{4}$) acting on (1,4) pair. The probability of obtaining $n$th outcome is
@@
$=$ $\frac{1}{4}\mbox{Tr}(\Pi_{n}).$
Using Eqs.(1) and (2) in (3) pbej1 , one can get
$|\Phi_{n}\rangle=\frac{1}{2\sqrt{p_{n}}}\sum_{\alpha=1}^{4}\sqrt{\pi_{n\alpha}}|\phi_{n\alpha}\rangle_{23}|\phi_{n\alpha}^{*}\rangle_{14},$ (5)
where $|\phi^{*}\rangle$ denotes the complex conjugate of $|\phi\rangle$ in the computational basis. Note that after the measurement, the total system state (5) remains pure.
Now after tracing out (2,3) qubits the post-measurement state between Alice and Charlie for $n$th outcome becomes
@@
$=$ $\frac{\Pi_{n}^{*}}{\mbox{Tr}(\Pi_{n})},$
where $\Pi_{n}^{*}=\sum_{\alpha=1}^{4}\pi_{n\alpha}|\phi_{n\alpha}^{*}\rangle\langle\phi_{n\alpha}^{*}|$. We also determine the post-measurement state of (1,2) and (3,4) pairs; for details, see Appendix A. Equation (6) illustrates how Bob’s measurement fully determines the state of the pair (1, 4). This equation also highlights the direct relationship between the post-measurement state of the (1,4) pair and the POVM measurement element. Conventional thought suggests that entanglement at the (1,4) pair requires an entangled measurement. However, we demonstrate that although an entangled POVM is sufficient to obtain an entangled state at (1,4), it is not a necessary condition
@@
Now, we consider that Bob post-processes each outcome of the first measurement $\mathcal{M}_{1}$ and then performs a second measurement $\mathcal{M}_{2}$ on the (2,3) qubits of the total system state (5). Additionally, Bob classically communicates the outcomes to Alice and Charlie during the measurement process.
@@
$E_{m}=\sum_{\delta=1}^{4}\mu_{m\delta}|\psi_{m\delta}\rangle\langle\psi_{m\delta}|,$ (7)
where $|\psi_{m\delta}\rangle$, $\delta=1,2,3,4$, form a complete orthonormal basis on $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, $\mu_{m\delta}\geq 0$ for $\delta=1,2,3,4$.
The post-measurement state of the four-qubit system, when Bob’s outcomes for the successive measurements are the $n$-th element of the measurement $\mathcal{M}_{1}$ and the $m$th element of $\mathcal{M}_{2}$, is given by
@@
where it is understood that $E_{m}$ is acting on (2,3) pair and $\mathbb{I}_{4}$ is acting on (1,4) pair of $|\Phi_{n}\rangle$ which is given by Eq.(5). The probability of obtaining the $|\Phi_{nm}\rangle$ as the outcome state is $p_{n}s_{nm}$.
Here, we denote by $s_{nm}$ the probability of obtaining the $m$-th outcome of the second measurement, $\mathcal{M}_{2}$ when it acts on the $n$-th outcome of the first measurement $\mathcal{M}_{1}$. Using Eqs.(4), (5), and (7), we get
@@
Now using Eqs.(5), and (7), the above Eq.(8) can be written as
$|\Phi_{nm}\rangle$ $=$ $\frac{1}{2\sqrt{p_{n}s_{nm}}}\sum_{\alpha,\delta=1}^{4}\sqrt{\pi_{n\alpha}}\sqrt{\mu_{m\delta}}\langle\psi_{m\delta}|\phi_{n\alpha}\rangle|\psi_{m\delta}\rangle_{23}\otimes|\phi_{n\alpha}\rangle_{14},$
where $p_{n}$ and $s_{nm}$ are given by Eqs.(4) and (9) respectively.
After tracing out the (2,3) pair from the above Eq.(II), we obtain the post-measurement state between A and C for each outcome, given by
$\rho_{14|nm}=\frac{1}{4p_{n}s_{nm}}\sum_{\alpha,\beta=1}^{4}\sum_{\delta=1}^{4}\mu_{m\delta}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\langle\phi_{n\alpha}|\psi_{m\delta}\rangle\langle\psi_{m\delta}|\phi_{n\beta}\rangle|\phi_{n\alpha}\rangle\langle\phi_{n\beta}|,$ (11)
where $|\phi_{n\alpha(\beta)}\rangle$, $\alpha(\beta)=1,2,3,4$, are orthonormal basis of the $n$th element $\Pi_{n}$ of the first measurement $\mathcal{M}_{1}$.
@@
Lemma 1. If the first measurement $\mathcal{M}_{1}$ is a projective or rank-1 POVM (pure POVM) measurement, meaning each $\Pi_{n}$ element is rank 1 or projective, then any second measurement $\mathcal{M}_{2}$ cannot disturb the post-measurement state of $\mathcal{M}_{1}$ between Alice and Charlie, whether by entangling, disentangling, or increasing the entanglement.
Proof: A POVM, $\mathcal{M}$ is called as rank 1 when all its elements are rank-1 operators. Suppose $\mathcal{M}_{1}$ is a projective or rank-1 POVM, implying each element $\Pi_{n}$ of $\mathcal{M}_{1}$ measurement is rank-1, such that for $\alpha=1$, $\Pi_{n}=\pi_{n1}|\phi_{n1}\rangle\langle\phi_{n1}|$. Then, according to Eq.5, $|\Phi_{n}\rangle$ can be expressed as:
@@
Thus far, we establish that for the second measurement $\mathcal{M}_{2}$ to disturb the (1,4) pair in the post-measurement state $|\Phi_{n}\rangle$, two conditions must be met: Firstly, the rank of the first measurement $\mathcal{M}_{1}$ should exceed one, and secondly, $|\Phi_{n}\rangle$ should exhibit non-zero entanglement in the $14|23$ bipartition (Eq.14). This prompts the following question: What are the properties of measurements that can lead to entangled states in (1,4) through successive measurements if the first attempt fails? To address this inquiry, we present protocols for different categories of measurements in the entanglement swapping scenario
@@ II.2 Separable operations
where we assume $|\Psi\rangle_{12}=|\phi^{+}\rangle$, a Bell state. Eqquation (30) reveals that any rank-2 measurement element retains the entanglement in the (1,2) and (3,4) pairs. Therefore, in the second round of measurement, if Bob conducts an inseparable operation on the (2,3) pair after post-processing the first measurement $\mathcal{M}_{1}$, there’s a possibility of obtaining an entangled state in the (1,4) pair. However, it’s crucial to note that the second measurement should be an inseparable operation, ensuring that the condition $\mathcal{C}_{n|12vs34}>0$ for an outcome [as stated in Eq. 17] is met.
It is noteworthy that our protocols are not only applicable for two rounds of successive measurements, but one can proceed further for $n$ rounds by following the same procedure until the (1,4) pair becomes entangled. After each round of measurement, the entanglement between the $14|23$ bipartition of the four qubits post-measurement state must be non-zero, for an outcome, in order for the next round of measurement to disturb the (1,4) pair, after post-processing the previous round measurements. We already know that a measurement will be successful in establishing entanglement in the (1,4) pair if it can create entanglement between the $12|34$ bipartition of the total system state.
@@ III CONCLUSIONS
Our findings demonstrate that while an entangled measurement is sufficient, it is not mandatory for the measurement operators to be entangled to establish an entangled state between A and C. We introduce a sequential approach where Bob conducts an initial measurement, followed by another measurement if the initial one fails to establish the entanglement. This sequential strategy, in conjunction with appropriate post-processing after the initial measurement, facilitates the establishment of entanglement between A and C. Our investigation offers unique insights into the role of measurement operators in entanglement generation.
@@
Future research efforts should prioritize refining entanglement activation protocols, exploring innovative measurement strategies, and broadening the protocol’s applicability to diverse scenarios, including nonlocality activation klobus ; gisin and network nonlocality bran1 ; bran2 . By addressing these research directions, we can deepen our understanding of entanglement generation and distribution, leading to advancements in quantum communication and computation.
@@ Acknowledgements
The authors express gratitude to Professor S. Bandyopadhyay for fruitful discussions. PB acknowledges S. Halder for helpful discussions.